When a 360-nF air capacitor is connected to a power supply, the energy stored in the

capacitor is 1.85 π‘₯10

βˆ’5

𝐽. While the capacitor is kept connected to the power supply, a slab

of dielectric is inserted that completely fills the space between the plates. This increases

the stored energy by 2.32 π‘₯10

βˆ’5

𝐽. (a) What is the potential difference between the

capacitor plates? (b) What is the dielectric constant of the slab?


1
Expert's answer
2021-05-04T12:06:26-0400

(a) Energy stored in capacitor without a dielectric, "U_o=1.85\times10^{-5}\space J"

Capacitance of capacitor, "C_o=360\space nF=360\times10^{-9}\space F"

Energy stored in a capacitor is given by

"U=\dfrac{1}{2}CV^2"

"V=\sqrt\dfrac{2U}{C}"

"V_o=\sqrt\dfrac{2\times1.85\times10^{-5}}{360\times10^{-9}}"

"V_o=10.137\space V"


(b) Energy stored in capacitor increases by "2.32\times10^{-5}\space J"

"U=U_o+(2.32\times10^{-5})"

"U=4.17\times10^{-5}\space J"


Energy stored in a capacitor increases when a dielectric slab is inserted as

"U=\dfrac{1}{2}kCV_o^2"

"k=\dfrac{2U}{CV_o^2}=\dfrac{2(4.17\times10^{-5})}{(360\times10^{-9})(10.137)^2}"

"k=2.254"